OpenStudy (anonymous):

a^(3-x)*b^(5x)=a^(3x)*b^(x+5) a^(3-x)*b^(5x)=a^(3x)*b^(x+5) @Mathematics

OpenStudy (agreene):

What's the question?

myininaya (myininaya):

for this to be equal we need 3-x=3x and 5x=x+5 3=4x 4x=5 3/4=x x=5/4 but x cannot be two different values at the same time

myininaya (myininaya):

no solution

OpenStudy (agreene):

No solution iff: a and be arent both 0, 1 or -1

myininaya (myininaya):

true

myininaya (myininaya):

is that what you wanted mirzahasnain?